If n = 6, the correct sequence of filling of electrons will be
(1) ns → np(n -1)d → (n - 2)f
(2) ns → n(n - 2)f → (n -1)d → np
(3) ns → (n -1)d → (n - 2)f → np
(4) ns → (n - 2)f → np → (n -1)d

Grade:upto college level

3 Answers

Navjyot Kalra
9 years ago
Sol: Fact Answer : (2)
Nitesh Singh Sagar
6 years ago
Aufban principle (n+l rule):Electron will fill orbit by sum of quantum numbers n and l. Orbitals with equal values of n+l will fill with the lower n values first.Since n=6 :ns⇒6s, n+l=6+0=6ns⇒6s,n+l=6+0=6np⇒6p, n+l=6+1=7np⇒6p,n+l=6+1=7(n−1)d⇒5d, n+l=5+2=7(n−2)f⇒4f, n+l=4+3=7Since in last three value of n+l value is same. Therefore we will with the lower n values first.So correct order,ns→(n−2)f→(n−1)d→np
Kushagra Madhukar
3 years ago
Dear student,
 
As per the n + l rule, we have n = 6
and also realising the fact that for same n + l value the orbital with lower n value will be filled first.
Hence our answer will be
6s → 4f → 5d → 6p
i.e. ns → (n – 2)f → (n – 1)d → np
Hence, option (2) is correct.
 
Thanks and regards,
Kushagra